# for Data-Driven Science

### v2

@@ Abstract
Ken Gu1 Ruoxi Shang1 Ruien Jiang2* Keying Kuang2* Richard-John Lin3* Donghe Lyu4* Yue Mao5* Youran Pan3* Teng Wu6* Jiaqian Yu7* Yikun Zhang1* Tianmai M. Zhang1* Lanyi Zhu1* Mike A. Merrill1 Jeffrey Heer1 Tim Althoff1 1University of Washington 2UC Berkeley 3New York University 4Stanford University 5University of British Columbia 6Microsoft 7George Washington University  https://github.com/behavioral-data/BLADE
@@ 1 Introduction
Scientific data continues to accumulate rapidly, driven by advancements in scientific instrumentation and the digitization of information. However, practicing data-driven science (i.e., answering research questions from data) remains difficult, requiring rigorous methodologies, an understanding of data values and semantics, statistical and domain expertise, and critical thinking to validate hypotheses and draw meaningful and justifiable conclusions Jun et al. (2021); Breznau et al. (2022); Baker (2016); Collaboration (2015).
@@
Language model (LM)-based agents Sumers et al. (2023); Wu et al. (2023); Wang et al. (2023), pre-trained on web-scale data and equipped with memory and tool usage capabilities Schick et al. (2023), have the potential to conduct and support data-driven science. They can reason about and interact with heterogeneous data representing subjects, objects, and processes of study in the “external” world Majumder et al. (2024b). However, to facilitate their progress, we need a reliable method to evaluate and measure their performance.
@@
Recent benchmarks have enabled progress. However, they focus on either (1) data analysis execution with straightforward tasks containing a single, final, easily evaluated answer (e.g., Calculate the mean and standard deviation of the "Mar.2019" column Hu et al. (2024a); Yin et al. (2022); Liu et al. (2024a)) or (2) tasks for machine learning (ML) (e.g., improve the accuracy of an ML model Hong et al. (2024a); Huang et al. (2023b); Guo et al. (2024b)). For scientific analyses, these tasks require limited integration of external knowledge, limited understanding of data semantics, and limited grounding in external scientific knowledge. In addition, these benchmarks evaluate only on single metrics, such as ML model accuracy or completion rate. However, in the process of data-driven scientific discovery, the many intermediary decisions in a multi-step analysis are themselves critical to identify, meaningfully assess, and differentiate in order to improve agent performance.
@@
Evaluating agent performance on open-ended data-driven analyses, especially automatically, poses specific challenges. First, the natural flexibility in making analysis decisions Gelman and Loken (2014, 2019); Simmons et al. (2011) makes it difficult to establish a single ground truth that encompasses all justifiable choices. Second, the heterogeneity of decisions (e.g., regarding specific hyperparameters of a statistical model, choices of variables, high-level approaches, etc.) complicates efforts to decide on the representation and abstraction of meaningful decisions. Finally, given multiple valid decisions and approaches, determining the criteria and method to assess the correctness and soundness of the agent’s analysis is difficult to quantify.
@@
In this work, we introduce BLADE, a benchmark for the principled evaluation of LM agents used for data-driven scientific analyses. Given a research question (e.g., “Are soccer players with a dark skin tone more likely than those with a light skin tone to receive red cards from referees?” Silberzahn et al. (2018); Auspurg and Brüderl (2021)) and a dataset, BLADE evaluates agents’ ability to integrate external scientific and statistical knowledge with an understanding of the data to conduct rigorously justifiable data analyses.
@@
To build BLADE, we collected a set of actual research questions and datasets (Fig. 1.1) from research papers, crowd-sourced analysis studies, and statistics textbooks (Sec. 3). Then, inspired by prior crowd-sourced analysis studies Silberzahn et al. (2018); Schweinsberg et al. (2021), we recruited expert data analysts and collected high-quality data analyses (Fig. 1.2) through a crowd-sourced analysis for each research question (i.e., multiple analysts independently performing a single analysis). To ensure our benchmark captured a broad variety of defensible analysis approaches, we asked analysts to validate alternative decisions from their peers and LM-generated decisions seeded by analysts’ own decisions. For this process, we also collected negative examples of "unjustifiable" decisions to use when testing agents’ ability to discern justifiable ones. We then combined all unique decisions to form the ground truth (Fig. 1.3).
@@
Next, based on studies outlining decision steps in the data analysis process Gu et al. (2023a); Liu et al. (2019, 2020a); Jun et al. (2021), we formulated tasks. These tasks tested the discernment and formulation of analytical decisions that reflect multiple levels of abstraction, ranging from executable code implementing data transformations to higher-level planning of conceptual variables requiring external scientific knowledge (Sec. 4).
@@
Hong et al. (2024b) Huang et al. (2023a) Liu et al. (2024b) Guo et al. (2024b) Hu et al. (2024b)
@@
(1) comprehend data semantics -- -- -- --
(2) integrate domain knowledge -- -- -- --
(3) conduct multi-step reasoning -- --
@@
(7) input-flexible decisions
Table 1: Comparing BLADE against existing data analysis evaluation datasets and benchmarks for conducting scientific analyses based on the requirements specified in Section 2. -- indicates partial satisfaction (e.g., data understanding is only on ML model building). See Table 4 for examples from BLADE and recent benchmarks.
Overall, BLADE contains 188 multiple choice and 536 ground truth analysis decisions encompassing multiple justifiable analysis approaches across 12 real-world datasets and research questions. To illustrate its utility and assess benchmark performance, we evaluate different LMs and a standard ReAct agent Yao et al. (2023) that interacts with a sandbox notebook environment (Sec. 6).
@@
Our work takes the first step in evaluating the use of agents for open-ended data-driven scientific discovery. It advances our understanding of agent capabilities to ultimately collect datasets, generate hypotheses, conduct analyses, and interpret results to form valid and justifiable scientific conclusions. To facilitate further research and development, we have made our benchmark and evaluation framework publicly available [^1] https://github.com/behavioral-data/BLADE.
@@ 2 Benchmark Requirements
Our benchmark evaluates agents on answering open-ended data-driven scientific questions, advancing current efforts that execute analysis code based on precise single-answer instructions. Existing benchmarks are limited in their ability to assess agent decision-making during analysis and often do not capture the full scope of their approaches. As summarized in Table 1, our benchmark addresses these limitations by focusing on the following key requirements.
We maintain that the ideal benchmark would evaluate an agent’s abilities to (1) comprehend data semantics, understanding the semantic relationships between variables and what the data represents relative to the external world, (2) integrate domain knowledge, i.e., findings from related literature and an understanding of a “world model”, (3) conduct multi-step reasoning and planning at different levels of abstraction, i.e., high level planning vs. lower level code execution, given domain knowledge, an understanding of the data, and execution outputs, and (4) differentiate justifiable decisions with firm theoretical or statistical support Simonsohn et al. (2020) from unjustifiable ones.
Additionally, the benchmark evaluation should be (5) automatic, requiring no human intervention, (6) decision-based, with the ground truth reflecting the intermediary decisions, and (7) input flexible decisions, being aware of multiple ways to specify the same decision.
Requirements 1 through 4 inform our data collection process (Sec. 3) and task formulation (Sec. 4). Requirements 5 through 7 inform our evaluation procedure (Sec. 5). Ultimately, we assess whether agents can plan, develop, and execute a justifiable analysis to answer a real-world research question.
@@ 3 Benchmark Data Collection
We now describe our data collection process for research questions (RQs), datasets, and ground-truth analyses.
RQs and Data. We selected scientific-grade datasets and RQs directly from scientific publications, particularly those studied in meta-analysis papers Silberzahn et al. (2018); Simonsohn et al. (2020); Young and Holsteen (2017) and reproduced in statistics textbooks Mcelreath (2020); Kleiber and Zeileis (2008). We chose these sources because they provide a multitude of complex analyses, and relevant properties that make analyses non-trivial and revealing of statistical knowledge. Table 3 summarizes these RQs, datasets, source papers, and meta-analysis papers. During this process, we ensured that the datasets were clearly documented and were sufficiently complex to require non-trivial analyses, i.e., expert annotators would be required to clearly distinguish defensible from indefensible decisions.
Annotation Process. To gather ground truth analyses and ensure the highest quality annotations, we followed a procedure similar to those used in previous crowd-sourced analysis studies Silberzahn et al. (2018); Schweinsberg et al. (2021). We recruited 11 trained analysis experts and engaged them in a multi-stage process to ensure quality. Our experts had a self-reported average of 6 years of experience, with 6 pursuing or holding a Ph.D. in a scientific field. Since one of our key contributions is the corpus of ground truth analyses, we invited our expert annotators to be co-authors of this paper. See Appendix A.1 for details on analyst recruitment.
We gave each expert an RQ, dataset, and dataset description, including details of each column. For each dataset, experts independently conducted their analyses, recording all decisions they made. This naturally resulted in multiple analytical approaches. To broaden the scope of possible strategies, we used an LM (GPT-4) to generate additional decisions so that we could capture a broad diversity of alternative approaches (prompts shown in Fig. 9). The LM prompt was seeded with existing expert annotations. In addition, the LM-generated decisions were also used to gather labelled examples of unjustifiable decisions. Specifically, to ensure high data quality, experts validated and annotated each other’s and LM-generated decisions as justified or unjustified.
Agreement rates among expert annotators were relatively high: 75% for transformations and 80% for conceptual variables. In contrast, agreement on LM-generated decisions was much lower, at 27% for transformations and 13% for conceptual variables. As a result, fewer than one-third of the LM-generated transformations and conceptual variables were unanimously approved, and many of these lower-agreement items were excluded from the final ground truth, highlighting areas where LM agents need improvement.
@@ 4 Benchmark Tasks
We want the benchmark tasks to represent decisions that are vital to the analysis and to evaluate the key skills needed to conduct data-driven science (i.e., requirements 1-4 in Section 2). Below, we introduce the decisions BLADE tests that reflect these skills before we examine specific BLADE tasks.
4.1 Types of Decisions Tested
Drawing from prior work studying the scientific analysis process Gu et al. (2023b); Liu et al. (2019, 2020a), we focus on an agent’s ability to make planning decisions, i.e., those requiring a process of reasoning about and then synthesizing the data, scientific domain, and statistical knowledge. In addition, we extend prior benchmarks that exclusively focused on the execution of data analysis by evaluating analysis execution in the context of a research question and higher-level analysis plans.
Specifically, BLADE measures the following decisions that encapsulate a data analysis.
  1. 1.
    Formulating Conceptual Variables. This involves identifying the high-level variables in an analysis and how they are tied to external information, e.g., "Prior literature suggests player physicality influences the referee’s perception of the player." Here, integrating scientific domain knowledge and conducting multi-step reasoning (requirements 2 and 3) are at issue. In the context of a RQ and analysis, this means to make these decisions, the agent must be able to identify the constructs that would serve as the independent variables (IVs), dependent variable (DV), and any control variables.
  2. 2.
    Executing Data Transformations. Given the research question and dataset, agents must be able to identify the relevant columns and transforms to the data to operationalize conceptual variables, e.g., be familiar enough with the semantics of the data to know that body mass index (BMI) could be a suitable proxy for player physicality using the “weight” and “height” columns.
  3. 3.
    Implementing Statistical Models. These decisions involve implementing the appropriate statistical model given the conceptual variables and transformed data to answer the research question. Doing so requires an in-depth knowledge of both statistical methods and the underlying conceptual hypothesis Jun et al. (2019, 2021).
@@ 4.1 Types of Decisions Tested
4.2 Specific Tasks Tested
Task 1: Discern Justifiable Decisions. To evaluate how well agents can discern justifiable decisions (requirement 4), BLADE includes the following Multiple Choice Questions. (MCQ1) Given the research question and the dataset, which conceptual variable is the most/least justifiable for the analysis? (MCQ2) Given the research question, dataset, and conceptual variable of interest, which transformation is the most/least justifiable to operationalize the variable?
@@ 4.2 Specific Tasks Tested
Task 2: Generate an End-to-end Analysis. For this significantly more complex task, agents need to generate a complete end-to-end analysis given a research question and a dataset. Specifically, to test agent performance on key analysis decisions (Sec. 4.1), agents are to submit the following artifacts (e.g., in Fig. 1.5 and 7), each mapping to one type of decision.
@@ 5 Flexible Automatic Evaluation
To quantitatively measure the quality of agent-generated analyses (i.e., the agent-generated artifacts) in a way that is automatic, decision-based, and input flexible (requirements 5-7 in Sec. 2), we need both concrete representations of analysis decisions and associated matching criteria. We now discuss the representation and matching procedure for each artifact in an agent’s submission.
@@
Matching Conceptual Variables. Because conceptual variables capture high-level constructs, two similarly specified constructs (i.e, player physicality and how physically imposing the player is) should have the same meaning as long as they have the same variable type (i.e., IV, DV, or Control). To match these specifications, we employed an LM (GPT-4o) to determine the semantic equivalence between two conceptual variables. We followed a procedure similar to  Liang et al. (2023) which was validated on semantically matching academic reviews (prompt in Fig. 11). Appendix A.4.2 contains further details.
@@
Matching Data Transformations. Since there are many ways to express data analyses in code, even ways that could be perfectly equivalent, we require a representation that maps equivalent transformations to a single representation (i.e., requirement 7 – input flexible evaluation). Taking the code in Figure 2 as an example, the ordering of the transforms 1- 2- 3 or 2- 1- 3 are functionally equivalent with respect to the final product of the computation (i.e., as long as 1 and 2 come before 3). In addition, transformations that result in the exact same output column values (with a small margin of error for floating point) should be considered equivalent transformations. Likewise, getting a certain column’s values correct should mean that all relevant prior steps were correct and that decisions for each relevant prior transformation were correct. For example, if a submission missed 6 but still correctly calculated the “rdcards” after the groupby, then the agent still correctly performed steps 1- 5, deserving significant partial credit. In complex tasks such as scientific data analysis, such partial credit enables meaningful differentiation of model performance and progress.
@@
To capture the aforementioned nuances, we developed a representation for data transformations using a data flow graph Kavi et al. (1986) (Fig. 2 right). These graphs are useful because any series of transformations in the order of a topological sort Manber (1989) on the graph leads to the same result. In addition, our graph captures data flow at the column-level (i.e., all cell values in a single column) to enable subsequent matching at the granularity of columns. In doing so, we allow for matching on transforms that require and affect only a subset of columns in a data table (e.g., in Fig. 2, getting 1 correct is independent of getting 2 correct). Appendix A.3 describes our data flow graph formalism in greater detail.
In addition, the transforms (i.e., orange nodes in Fig. 2) in the data flow graph represent a discrete data transformation decision that was made in wrangling the data (requirement 6 – decision-based evaluation). Specifically, each transform is defined by a fixed set of transform verbs (Table 6) that are: based on existing data wrangling libraries (i.e., Arquero [^2] https://idl.uw.edu/arquero/ and Vega Satyanarayan et al. (2016)), expandable, and validated to cover every analysis decision in our benchmark.
@@
To match transforms in BLADE, we applied an LM (GPT-4o) to convert the transformation function in an agent’s submission to the individual transform units (prompt in Fig. 12 and  13). We then constructed the agent’s transformation data flow graph and matched it with the ground truth. We match based on both the column values that are the output of any discrete transformation and a fuzzier graph isomorphism matching that determines whether approximately the same steps were applied. Appendix A.4.1 describes the matching procedures in detail.
@@
Matching Statistical Models. The implementation of statistical models and relevant parameters could be evaluated in multiple different ways (i.e., code, natural language, or mathematical formulas  Jun et al. (2021); McElreath (2018)). To prioritize the underexplored planning aspects of statistical modeling Gu et al. (2023b), we focus on being able to select the right model and conceptual variables. In principle, this representation could be extended to include code, hyperparameters, and more. Specifically, we first used an LM (GPT-4o) prompt (Fig. 14) to convert the modeling function into a natural language specification of the model and the columns in the transformed data table that are used in modeling. Next, using another LM (GPT-4o) prompt, we compared this output with the ground truth natural language specifications of the model (prompt in Fig. 15) and associated conceptual variables based on semantic equivalence. See Appendix A.4.3 for additional details.

### v3

@@ Abstract
Ken Gu1 Ruoxi Shang1 Ruien Jiang211footnotemark: 1 Keying Kuang211footnotemark: 1 Richard-John Lin311footnotemark: 1 Donghe Lyu411footnotemark: 1 Yue Mao511footnotemark: 1 Youran Pan311footnotemark: 1 Teng Wu611footnotemark: 1 Jiaqian Yu711footnotemark: 1 Yikun Zhang [^1] footnotemark: 1 Tianmai M. Zhang [^1] footnotemark: 1 Lanyi Zhu1thanks: These authors contributed equally to this work. Mike A. Merrill1 Jeffrey Heer1 Tim Althoff1 1University of Washington 2UC Berkeley 3New York University 4Stanford University 5University of British Columbia 6Microsoft 7George Washington University  https://github.com/behavioral-data/BLADE
@@ 1 Introduction
Scientific data continues to accumulate rapidly, driven by advancements in scientific instrumentation and the digitization of information. However, practicing data-driven science (i.e., answering research questions from data) remains difficult, requiring rigorous methodologies, an understanding of data values and semantics, statistical and domain expertise, and critical thinking to validate hypotheses and draw meaningful and justifiable conclusions Jun2021HypothesisFE; Breznau2022ObservingMR; Baker20161500SL; Aarts2015EstimatingTR.
@@
Language model (LM)-based agents Sumers2023CognitiveAF; Wu2023AutoGenEN; Wang2023ASO, pre-trained on web-scale data and equipped with memory and tool usage capabilities Schick2023ToolformerLM, have the potential to conduct and support data-driven science. They can reason about and interact with heterogeneous data representing subjects, objects, and processes of study in the “external” world Majumder2024DatadrivenDW. However, to facilitate their progress, we need a reliable method to evaluate and measure their performance.
@@
Recent benchmarks have enabled progress. However, they focus on either (1) data analysis execution with straightforward tasks containing a single, final, easily evaluated answer (e.g., Calculate the mean and standard deviation of the "Mar.2019" column Hu2024InfiAgentDABenchEA; Yin2022NaturalLT; Liu2024AreLC) or (2) tasks for machine learning (ML) (e.g., improve the accuracy of an ML model Hong2024DataIA; Huang2023MLAgentBenchEL; Guo2024DSAgentAD). For scientific analyses, these tasks require limited integration of external knowledge, limited understanding of data semantics, and limited grounding in external scientific knowledge. In addition, these benchmarks evaluate only on single metrics, such as ML model accuracy or completion rate. However, in the process of data-driven scientific discovery, the many intermediary decisions in a multi-step analysis are themselves critical to identify, meaningfully assess, and differentiate in order to improve agent performance.
@@
Evaluating agent performance on open-ended data-driven analyses, especially automatically, poses specific challenges. First, the natural flexibility in making analysis decisions Gelman2014TheSC; Gelman2019TheGO; Simmons2011FalsePositiveP makes it hard to establish a single ground truth that encompasses all justifiable choices. Second, the heterogeneity of decisions (e.g., regarding hyperparameters of a statistical model, variables choices, high-level approaches, etc.) complicates efforts to decide on the representation and abstraction of meaningful decisions. Finally, given multiple valid decisions and approaches, determining the criteria and method to assess the correctness and soundness of the agent’s analysis is difficult to quantify.
@@
In this work, we introduce BLADE, a benchmark for the principled evaluation of LM agents used for data-driven scientific analyses. Given a research question (e.g., “Are soccer players with a dark skin tone more likely than those with a light skin tone to receive red cards from referees?” Silberzahn2018ManyAO; auspurg2021has) and a dataset, BLADE evaluates agents’ ability to integrate external scientific and statistical knowledge with an understanding of the data to conduct rigorously justifiable data analyses.
@@
To build BLADE, we collected a set of actual research questions and datasets (Fig. 1.1) from research papers, crowd-sourced analysis studies, and statistics textbooks (Sec. 3). Then, inspired by prior crowd-sourced analysis studies Silberzahn2018ManyAO; Schweinsberg2021SameDD, we recruited expert data analysts and collected high-quality data analyses (Fig. 1.2) through a crowd-sourced analysis for each research question (i.e., multiple analysts independently performing a single analysis). To ensure our benchmark captured a broad variety of defensible analysis approaches, we asked analysts to validate alternative decisions from their peers and LM-generated decisions seeded by analysts’ own decisions. For this process, we also collected negative examples of "unjustifiable" decisions to use when testing agents’ ability to discern justifiable ones. We then combined all unique decisions to form the ground truth (Fig. 1.3).
@@
Next, based on studies outlining decision steps in the data analysis process Gu2023HowDD; Liu2019PathsEP; Liu2020UnderstandingTR; Jun2021HypothesisFE, we formulated tasks. These tasks tested the discernment and formulation of analytical decisions that reflect multiple levels of abstraction, ranging from executable code implementing data transformations to higher-level planning of conceptual variables requiring external scientific knowledge (Sec. 4).
@@
hong2024data huang2023benchmarking liu2024llms Guo2024DSAgentAD hu2024infiagent
@@
(1) comprehend data semantics - - - -
(2) integrate domain knowledge - - - -
(3) conduct multi-step reasoning - -
@@
(7) flexible to decision input
Table 1: Comparing BLADE against existing data analysis evaluation datasets and benchmarks for conducting scientific analyses based on the requirements specified in Section 2. - indicates partial satisfaction (e.g., data understanding is only on ML model building). See Table 4 for examples from BLADE and recent benchmarks.
Overall, BLADE contains 188 multiple choice and 536 ground truth analysis decisions encompassing multiple justifiable analysis approaches across 12 real-world datasets and research questions. To illustrate its utility and assess benchmark performance, we evaluate different LMs and a standard ReAct agent yao2023react that interacts with a sandbox notebook environment (Sec. 6).
@@
Our work takes the first step in evaluating agents for open-ended data-driven scientific discovery, advancing our understanding of their capabilities to collect data, generate hypotheses, conduct analyses, and interpret results to form valid, justifiable scientific conclusions. To support further research and development, we open source our benchmark and evaluation framework [^1] https://github.com/behavioral-data/BLADE.
@@ 2 Benchmark Requirements
Our benchmark evaluates agents on answering open-ended, data-driven scientific questions, advancing current efforts that execute analysis code based on precise single-answer instructions. Existing benchmarks are limited in their ability to assess agent decision-making during analysis and often do not capture the full scope of their approaches. Our benchmark addresses these limitations by focusing on the following key requirements (Table 1).
We maintain that the ideal benchmark would evaluate an agent’s abilities to (1) comprehend data semantics, understanding the semantic relationships between variables and what the data represents relative to the external world, (2) integrate domain knowledge, i.e., findings from related literature and an understanding of a “world model”, (3) conduct multi-step reasoning and planning at different levels of abstraction, i.e., high level planning vs. lower level code execution, given domain knowledge, an understanding of the data, and execution outputs, and (4) differentiate justifiable decisions with firm theoretical or statistical support simonsohn2020specification from unjustifiable ones.
Additionally, the benchmark evaluation should be (5) automatic, requiring no human intervention, (6) decision-based, with the ground truth reflecting the intermediary decisions, and (7) flexible to decision input, being aware of multiple ways to specify the same decision. Requirements 1 through 4 inform our data collection process (Sec. 3) and task formulation (Sec. 4). Requirements 5 through 7 inform our evaluation procedure (Sec. 5). Ultimately, we assess whether agents can plan, develop, and execute a justifiable analysis to answer a real-world research question.
@@ 3 Benchmark Data Collection
We now describe our data collection process for research questions (RQs), data, and ground-truth analyses.
RQs and Data. We selected scientific-grade datasets and RQs directly from scientific publications, particularly those studied in meta-analysis papers Silberzahn2018ManyAO; simonsohn2020specification; young2017model and reproduced in statistics textbooks Mcelreath2020StatisticalR; kleiber2008applied. We chose these sources because they provide a multitude of complex analyses, and relevant properties that make analyses non-trivial and revealing of statistical knowledge. Table 3 summarizes these RQs, datasets, source papers, and meta-analysis papers. During this process, we ensured that the datasets were clearly documented and sufficiently complex to require non-trivial analyses, i.e., expert annotators would be required to distinguish defensible from indefensible decisions.
Annotation Process. To gather ground truth analyses and ensure the highest quality annotations, we followed a procedure similar to those used in previous crowd-sourced analysis studies Silberzahn2018ManyAO; Schweinsberg2021SameDD. We recruited 11 trained analysis experts and engaged them in a multi-stage process to ensure quality. Our experts had a self-reported average of 6 years of experience, with 6 pursuing or holding a Ph.D. in a scientific field. Since one of our key contributions is the corpus of ground truth analyses, we invited our expert annotators to be co-authors of this paper. See Appendix A.1 for details on recruitment.
We gave each expert an RQ, dataset, and dataset description, including details of each column. For each dataset, experts independently conducted their analyses, recording all decisions they made. This naturally resulted in multiple analytical approaches. To broaden the scope of possible strategies, we used an LM (GPT-4) to generate additional decisions (prompts shown in Fig. 9).
To ensure high data quality, experts validated and annotated each other’s and LM-generated decisions as justified or unjustified. Agreement rates among expert annotators were relatively high: 75% for transformations and 80% for conceptual variables. In contrast, agreement on LM-generated decisions was much lower, at 27% for transformations and 13% for conceptual variables, highlighting low agent performance on decision generation. Many of these lower-agreement decisions were excluded from the final ground truth dataset.
@@ 4 Benchmark Tasks
We want the benchmark tasks to represent decisions that are vital to the analysis and to evaluate the key skills needed to conduct data-driven science (i.e., requirements 1-4 in Section 2). We draw from prior studying the scientific analysis process Gu2023HowDA; Liu2019PathsEP; Liu2020UnderstandingTR and focus on agents’ ability to discern (Sec. 4.1) and make (Sec. 4.2planning decisions, i.e., those requiring a process of reasoning about and then synthesizing the data, scientific domain, and statistical knowledge. Specifically we test the following decisions:
(1) Formulating Conceptual Variables. Agents should recognize independent variables (IVs) , dependent variables (DVs), and control variables based on domain knowledge and multi-step reasoning (requirements 2 and 3), e.g., “Prior literature suggests player physicality influences the referee’s perception. We can consider physicality a control.”
(2) Executing Data Transformations. Agents should select relevant columns and apply transformations to operationalize conceptual variables, e.g., using BMI as a proxy for player physicality via “weight” and “height” columns.
(3) Implementing Statistical Models. Agents should choose the appropriate statistical model based on conceptual variables and transformed data to address the research question, requiring in-depth knowledge of statistical methods and the underlying hypothesis Jun2019TeaAH; Jun2021HypothesisFE.
@@
4.1 Task 1: Discern Justifiable Decisions
To evaluate how well agents can discern justifiable decisions (requirement 4), BLADE includes the following Multiple Choice Questions. (MCQ1) Given the research question and the dataset, which conceptual variable is the most/least justifiable for the analysis? (MCQ2) Given the research question, dataset, and conceptual variable of interest, which transformation is the most/least justifiable to operationalize the variable?
@@ 4.1 Task 1: Discern Justifiable Decisions
4.2 Task 2: Generate an End-to-end Analysis
For this significantly more complex task, agents need to generate a complete end-to-end analysis given a research question and a dataset. Specifically, to test agent performance on key analysis decisions, agents are to submit the following artifacts (e.g., in Fig. 1.5 and 7), each mapping to one type of decision.
@@ 5 Flexible Automatic Evaluation
To quantitatively measure the quality of agent-generated analyses (i.e., the agent-generated artifacts) in a way that is automatic, decision-based, and flexible to decision input (requirements 5-7 in Sec. 2), we need concrete representations of analysis decisions and associated matching criteria. We now discuss the representation and matching procedure for each artifact in an agent’s submission.
@@
Matching Conceptual Variables. Because conceptual variables capture high-level constructs, two similarly specified constructs (i.e, player physicality and how physically imposing the player is) should have the same meaning as long as they have the same variable type (i.e., IV, DV, or Control). To match these specifications, we employed an LM (GPT-4o) to determine the semantic equivalence between two conceptual variables. We followed a procedure similar to  Liang2023CanLL which was validated on semantically matching academic reviews (prompt in Fig. 11). Appendix A.4.2 contains further details.
@@
Matching Data Transformations. Since there are many ways to express data analyses in code, even ways that could be perfectly equivalent, we require a representation that maps equivalent transformations to a single representation (i.e., requirement 7 – flexible to decision input). Taking the code in Figure 2 as an example, the ordering of the transforms 1- 2- 3 or 2- 1- 3 are functionally equivalent with respect to the final product of the computation (i.e., as long as 1 and 2 come before 3). In addition, transformations that result in the exact same output column values (with a small margin of error for floating point) should be considered equivalent transformations. Likewise, getting a certain column’s values correct should mean that all relevant prior steps were correct and that decisions for each relevant prior transformation were correct. For example, if a submission missed 6 but still correctly calculated the “rdcards” after the groupby, then the agent still correctly performed steps 1- 5, deserving significant partial credit. In complex tasks such as scientific data analysis, such partial credit enables meaningful differentiation of model performance and progress.
@@
To capture the aforementioned nuances, we developed a representation for data transformations using a data flow graph Kavi1986AFD (Fig. 2 right). These graphs are useful because any series of transformations in the order of a topological sort Manber1989IntroductionTA on the graph leads to the same result. In addition, our graph captures data flow at the column-level (i.e., all cell values in a single column) to enable subsequent matching at the granularity of columns. In doing so, we allow for matching on transforms that require and affect only a subset of columns in a data table (e.g., in Fig. 2, getting 1 correct is independent of getting 2 correct). Appendix A.3 describes our data flow graph formalism in greater detail.
@@
In addition, the transforms (i.e., orange nodes in Fig. 2) in the data flow graph represent a discrete data transformation decision that was made in wrangling the data (requirement 6 – decision-based evaluation). Specifically, each transform is defined by a fixed set of transform verbs (Table 6) that are based on existing data wrangling libraries (i.e., Arquero [^2] https://idl.uw.edu/arquero/ and Vega satyanarayan2016vega), expandable, and validated to cover every analysis decision in our benchmark. To match transforms in BLADE, we applied an LM (GPT-4o) to convert the transformation function in an agent’s submission to the individual transform units (prompt in Fig. 12 and  13). We then constructed the agent’s transformation data flow graph and matched it with the ground truth. We match based on both the column values that are the output of any discrete transformation and a fuzzier graph isomorphism matching that determines whether approximately the same steps were applied. Appendix A.4.1 describes the matching procedures in detail.
@@
Matching Statistical Models. The implementation of statistical models and relevant parameters could be evaluated in multiple ways (i.e., code, natural language, or mathematical formulas  Jun2021HypothesisFE; mcelreath2018statistical). To prioritize the underexplored planning aspects of statistical modeling Gu2023HowDA, we focus on being able to select the right model and conceptual variables. In principle, this representation could be extended to include code, hyperparameters, and more. We first used an LM (GPT-4o) prompt (Fig. 14) to convert the modeling function into a natural language specification of the model and the columns in the transformed data table that it used. Next, using another LM (GPT-4o) prompt (Fig. 15), we compared this output with the ground truth natural language specifications of the model and associated conceptual variables based on semantic equivalence. See Appendix A.4.3 for additional details.